OpenStudy (ttop0816):

help me solve step by step please! ): If 3cot^2x - 5cotx = -2, then cotx = _____. -2, -1/3 -1, -2/3 1, 2/3 2, 1/3

OpenStudy (amorfide):

this is a quadratic, move the constant to the left, then use the quadratic formula, and if you dont know how to use cot on your calculator use 1/tanx or cosx/sinx

OpenStudy (ttop0816):

so then it would be 3cot^2 - 5cotx +2=0

OpenStudy (ttop0816):

i would factor it and it would become (3cotx -1) (5cotx-2) =0??

OpenStudy (ttop0816):

im not quite sure about the factoring part;

OpenStudy (amorfide):

it would be easier for you to use the formula lol

OpenStudy (amorfide):

if you really wanted to factorise it do it step by step like

OpenStudy (ttop0816):

is it this formula? https://media.glynlyon.com/g_trg_2013/2/quad_formula.gif

OpenStudy (amorfide):

3cot^2x - 5cotx = -2 3cot^2x - 5cotx +2=0 3x2=6 factors of 6 1 6 2 3 what adds to give -5 but multiplies to give 6 -2 x -3 3cot^2x -3cotx -2cotx +2 factorise this 3cotx(cotx-1)-2(cotx-1) (3cotx-2)(cotx-1)

OpenStudy (amorfide):

yes that is the formula